In how many ways can 3 students be seated in a classroom with 5 chairs? Ans. 𝑃3 5 .
Found 2 solutions by saw, ikleyn:
Answer by saw(34)

nPr = 5!/(5-3)! = 5!/2! = 120/2 = 60
3 students can be seated with 5 chairs in 60 ways

Answer by ikleyn(52778)

.

The correct answer is 5*4*3 = 60.


-----------------

Hey, I do not understand precisely the meaning of your post.

For what reason you write in the post this fragment " Ans. P 3 5 " ?